OpenStudy (anonymous):

Which of the following is irrational? A.√10 B.√25 C.√100 D.√225

OpenStudy (anonymous):

a,

OpenStudy (anonymous):

A.√10

OpenStudy (across):

\[5^2=25\]\[10^2=100\]\[15^2=225\]\[3.1622776601683793319988935444327^2\approx10\]